# MLlm-DR: Towards Explainable Depression Recognition with MultiModal Large Language Models

### v1

@@ Abstract.
Automated depression diagnosis aims to analyze multimodal information from interview videos to predict participants’ depression scores. Previous studies often lack clear explanations of how these scores were determined, limiting their adoption in clinical practice. While the advent of LLMs provides a possible pathway for explainable depression diagnosis, current LLMs capable of processing multimodal data lack training on interview data, resulting in poor diagnostic performance when used directly. In this paper, we propose a novel multimodal large language model (MLlm-DR) that can understand multimodal information inputs and supports explainable depression diagnosis. MLlm-DR integrates a smaller LLMs and a lightweight query module (LQ-former). Specifically, the smaller LLMs is designed to generate depression scores and corresponding evaluation rationales. To enhance its logical reasoning for domain-specific tasks while maintaining practicality, we constructed a robust training dataset to fine-tune it. Meanwhile, the LQ-former captures depression-related features from speech and visual data, aiding the model’s ability to process multimodal information, to achieve comprehensive depression diagnosis. Our approach achieves state-of-the-art results on two interview-based benchmark datasets, CMDC and E-DAIC-WOZ, demonstrating its effectiveness and superiority.
@@ 1. Introduction
Interviews are widely regarded as the gold standard for diagnosing depression (gold, ). Clinicians evaluate the severity of symptoms by asking participants questions based on well-established diagnostic criteria, such as the DSM-5, ICD-11, and CCMD-3 (DSM-5, ; ICD-11, ; CCMD-3, ). These symptoms include suicidal tendencies, depressive mood, loss of interest, sleep disorders, among others. By combining the scores for each symptom, clinicians determine the final evaluation of depression. However, this process is often time-consuming and can be affected by the clinicians’ subjective judgment.
In recent years, significant progress has been made in automated depression diagnosis based on interviews (sub-attentional, ; MMFF, ; End-to-end, ). These studies analyze the semantics of dialogues and extract emotional cues from speech and facial expressions, leading to efficient and accurate diagnoses (iifdd, ; multimodal, ; MDDR, ; mago, ). However, a common limitation of these methods is their lack of explainability. While they predict participants’ depression scores using neural networks, they do not provide insights into how these scores are determined. This lack of clear rationale can lead to skepticism among clinicians, which restricts the broader adoption of these methods in clinical practice.
@@
Explainable depression diagnosis demands that diagnostic models not only output depression scores but also generate explanations for why such scores are given, thereby enhancing the credibility and acceptability of the diagnostic results. The advent of large language models (LLMs) (qwen2.5, ; llama, ; gpt4, ; ernie, ; glm, ) offers a promising pathway to achieving such explainable diagnostics. Specifically, LLMs excel in multi-turn dialogue tasks (zhaochat, ; Instructerc, ; fado, ; huang, ), enabling them to understand and analyze emotional, contextual, and semantic information within interview dialogue to assess participants’ depression scores. Moreover, their powerful logical reasoning (distilling, ; chain, ; kojima, ) and text generation abilities (T5, ; qwen2.5, ) allow them to generate detailed and coherent rationales based on the depression scores they assign. For example, if the model determines a high depression score, it can provide an explanation that highlights specific statements or content from the dialogue that correspond to depressive symptoms.
While LLMs are effective for understanding dialogue content in explainable depression diagnosis, the diagnostic indicators for depression are not limited to participants’ self-reported conversational content alone. They also include objective information, such as the patients’ voice and visual cues (tmac, ; facial, ). These factors are vital as they provide significant insights into an individual’s emotional and cognitive states, thereby offering a more comprehensive set of indicators for depression. Although some existing multimodal LLMs (2023gemini, ; 2024claude, ; Qwen2.5-VL, ) are capable of directly analyzing video data for zero-shot prediction, raw videos are often not provided in depression datasets due to patient privacy constraints, rendering these models inapplicable in real-world scenarios.
In this paper, we propose a novel multimodal large language model (MLlm-DR) that can understand multimodal inputs and supports explainable depression diagnosis, as shown in Figure  1. MLlm-DR is a strong practical framework that utilizes a smaller LLMs yet maintains exceptional logical reasoning capabilities in domain-specific tasks. To achieve this, we leverage advanced LLMs to generate evaluation rationales based on dialogue content, constructing a robust training dataset. This dataset is then used to fine-tune MLlm-DR, ensuring coherent and reliable evaluation rationales.
To further enhance the capabilities of MLlm-DR, enabling it to process multimodal information beyond text, we introduce a lightweight query module (LQ-former). LQ-former utilizes a set of learnable query vectors to extract depression-related features from speech and visual inputs, mapping these features into a unified text feature space compatible with the LLMs. These processed features are then fed into the frozen LLMs for text generation tasks, thereby training the LQ-former. After the training of LQ-former, we freeze it to preserve the learned feature extraction capability. Subsequently, we fine-tune the LLM using a joint optimization strategy that combines language modeling loss and regression loss, targeting both rationale generation and depression score prediction. This design not only improves the accuracy of depression recognition, but also provides clinically consistent and reliable assessments.
Our contributions can be summarized as follows:
  • We propose a novel MLlm-DR framework for explainable depression diagnosis. To the best of our knowledge, this is the first work to propose a multimodal LLMs specifically designed for depression recognition.
  • We construct an explainable depression assessment training data and fine-tune MLlm-DR, enhancing its logical reasoning capabilities while ensuring practicality.
  • We propose the LQ-former module, enabling LLMs to effectively integrate multimodal information for comprehensive depression diagnosis.
  • We evaluate our method on the CMDC and E-DAIC-WOZ datasets, achieving state-of-the-art results.
@@ 2.1. Multimodal Depression Recognition
Recently, multimodal fusion methods have made valuable progress in many depression recognition tasks. For instance, Wei et al. (sub-attentional, ) design independent attention fusion modules for each PHQ-8 sub-score to extract multi-modal features relevant to specific sub-scores, generating individual sub-scores and ultimately aggregating them for overall depression evaluation. Yuan et al. (MMFF, ) introduce multi-order factor decomposition to extract features from single modalities and their cross-modal combinations, significantly enhancing the representational capacity of multi-modal learning and improving model interpretability through a dynamic weighting mechanism. Jung et al. (HiQuE, ) explicitly model the hierarchical structure of interview questions (primary questions and follow-ups), simulating the diagnostic logic of clinicians while leveraging attention mechanisms to identify critical questions and features, enabling more precise depression detection. However, these methods rely on neural networks to directly predict depression scores without providing the corresponding rationale, which reveals a significant limitation.
@@ 2.2. LLMs-based Dialogue Understanding
Recently, large language models (LLMs) have made significant progress in dialogue understanding tasks. For instance, Li et al. (empirical, ) demonstrate the potential of LLMs in Dialogue Relation Extraction (DRE) tasks, showing their superior ability to capture long-span and multi-turn dialogue information, outperforming traditional sequence-based and graph-based methods. Lei et al. (Instructerc, ) investigate the use of LLMs for Emotion Recognition in Conversations (ERC). They pretrain the model on a speaker identification task to capture emotional expression characteristics of different roles in dialogues and fine-tune it through multi-task learning by integrating ERC and emotion influence prediction tasks, enhancing performance. Additionally, Huang et al. (huang, ) propose the Emotion-Cause Reasoning Chain (ECR-Chain) framework, leveraging LLMs to analyze statements in dialogues that trigger target emotions and thereby predict causal emotion entailment (CEE). This framework incorporates cognitive appraisal theory to deeply explore the process of emotion generation, thus providing strong interpretability for reasoning outcomes. These studies collectively demonstrate the tremendous potential of LLMs in advancing dialogue understanding tasks.
@@ 2.3. Cross-Modal Semantic Alignment
To enable LLMs to process and understand multimodal inputs, researchers have proposed various cross-modal semantic alignment methods(flamingo, ; cogvlm, ), achieving significant progress in multiple cross-modal language generation tasks, such as visual question answering and image captioning. For example, Li et al. (Q, ) propose a semantic alignment module (Q-Former), which utilizes a set of learnable query vectors to capture key information from visual features. Q-Former is trained through a vision-language matching task to align visual and textual features, enabling LLMs to generate textual descriptions based on visual input. To reduce reliance on paired image-text data, Jian et al. (P, ) training their model on textual data to optimize prompts that guide the generation of language from visual inputs. Visual features are subsequently mapped to these prompts, achieving alignment between vision and language. To enhance the model’s understanding of fine-grained visual information, Lu et al. (MQ, ) capture multi-level image features through three submodules: image tagging, object detection, and semantic segmentation. This enables the LLMs to process more granular visual information. Additionally, approaches (Emotion-LLaMA, ) employ simple linear networks (e.g., MLP) to map non-textual features into the textual embedding space, facilitating modality alignment. These studies provide effective methods to extend text-only LLMs to process multimodal data.
@@ 3.1. Data Processing
To enhance the performance of smaller LLMs in depression assessment tasks and enable them to generate more logically consistent evaluation rationales, we utilize advanced LLMs (gpt4, ; llama, ) to generate evaluation rationales based on dialogue content and construct training dataset. Specifically, we provide a text instruction requesting the advanced LLMs to produce an evaluation score within a range of 0-3 along with corresponding rationales based on the dialogue content. In this text instruction, we report the actual evaluation score of the patient, ensuring consistency between the rationales generated by the LLMs and the actual score, as shown in the Table 1. The training dataset is then used to fine-tune MLlm-DR, aiming to transfer the reasoning capabilities of the advanced LLMs to a smaller model for domain-specific tasks. In addition, we use HuBERT (hubert, ) to extract audio features from raw speech data. For visual features, we utilize the deep representations provided in the dataset, which are extracted using OpenFace 2.0 (OpenFace2.0, ) or ResNet 50 (ResNet, ).
@@
User Given the participant’s self-rating score of {label}, please evaluate the participant’s performance in {aspect} based on the dialogue content. The output format is as follows: Evaluation Result: A numeric value between 0-3. Evaluation Reason: A concise and logical description. Each output must strictly follow this format, avoiding omissions or confusion.
@@ 3.2. Lightweight Query Module
The LQ-former is designed to help LLMs process and understand multimodal information. It consists of a Transformer (attention, ) Decoder and a Projection Module based on a fully connected network. The Decoder takes a set of learnable query vectors $Q$ as input, which interacts internally via self-attention mechanisms and with non-textual modality features (such as audio and visual) via cross-attention mechanisms, thereby capturing depression cues within the non-textual modality features, and generating fixed-length feature representations $H$. $H$ are then the Projected to the same dimensionality as the LLM’s textual embeddings through a projection layer, producing the output $O$, as shown in the Equation  1. In our task, we utilize two LQ-former modules to process audio and visual features separately. The resulting features are concatenated with the tokenized textual instruction embeddings and used as inputs to the LLMs. To indicate the positions of the inserted features, we employ two special tokens, <AudioHere>and <VideoHere>, as markers.
@@
where $Q\in{\mathbb{R}^{n\times{D}}}$ is the query vector, while $K=V\in{\mathbb{R}^{m\times{D}}}$ represents the sequence of non-textual modality features, $D$ is the embedding dimension, $n$ denotes a fixed sequence length, and $m$ represents the length of the non-textual feature sequence,
@@ 3.3. Multi-Task Joint Fine-Tuning
Existing LLMs primarily focus on text generation tasks, and their performance in score prediction is often limited when the provided prompt lacks sufficient example samples. To enhance the model’s performance in scoring prediction tasks, we introduce a joint optimization strategy combining causal language modeling cross-entropy loss ($\mathcal{L}_{\text{lm}}$) and mean squared error loss ($\mathcal{L}_{\text{mse}}$), as shown in the Equation  2.
(2) $\mathcal{L}_{\text{joint}}=\mathcal{L}_{\text{lm}}+\mathcal{L}_{\text{mse}}.$
To process the regression prediction task, we design a Multi-Head Attention Aggregation Network (MAA). MAA takes the final hidden states $X\in\mathbb{R}^{L\times D}$ of the LLMs as input. It first splits $X$ into $h$ subspaces corresponding to $h$ attention heads, resulting in a tensor $x^{\prime}\in\mathbb{R}^{L\times h\times d}$, Each head then computes token-level attention weights $\alpha_{i}\in\mathbb{R}^{L}$ using a shared fully connected layer. These weights are used to aggregate the token representations into a sentence-level embedding. The embedding is passed through a fully connected layer to output the predicted scores $S_{\text{reg}}$, which are used to calculate the regression loss, as shown in the Equation  3. where $L$ is the sequence length, $h$ is the number of attention heads, and $d=D/h$ is the the dimension of each subspace, $W_{i}\in\mathbb{R}^{d\times 1}$ is a trainable parameter for each head.
$S_{\text{reg}}$ $=\text{FC}\left(\sum_{i=1}^{h}\left(\alpha_{i}\cdot x_{i}^{\prime}\right)\right).$

### v2

@@ Abstract.
Clinical depression diagnosis relies heavily on both verbal and non-verbal cues in patient interviews, yet existing automated methods often operate as black-box models and fail to provide trustworthy explanations, limiting their clinical applicability. Moreover, depression datasets commonly impose strict privacy constraints that prohibit access to raw audio–visual data, and the deployment of large LLMs in real-world medical environments is often constrained by computational cost limitations. To address these challenges, we propose MLlm-DR, a multimodal large language model for explainable depression recognition. We first employ knowledge distillation to transfer high-quality diagnostic rationales from a powerful LLMs to a smaller, deployable LLMs, thereby equipping it with clinically aligned reasoning abilities. We further introduce a lightweight query module (LQ-former) that extracts salient depression-related cues from pre-extracted audio and visual features and maps them into an LLM-compatible representation. MLlm-DR is trained in two stages: multimodal alignment via LQ-former pretraining, followed by multi-task optimization that jointly learns rationale generation and score regression, encouraging consistency and improving interpretability. Experimental results show that MLlm-DR achieves state-of-the-art performance on two interview-based benchmark datasets, CMDC and E-DAIC-WOZ, while simultaneously generating clinically meaningful and readable explanations.
@@ 1. Introduction
Clinical interviews are the cornerstone of depression diagnosis, widely regarded as the gold standard (Aalto et al., 2012). During these interviews, clinicians assess symptom severity based on established criteria like DSM-5, ICD-11, and CCMD-3 (American Psychiatric Association et al., 2013; Organization and others, 2018; of Psychiatry, 2001), evaluating indicators such as suicidal ideation, anhedonia, and sleep disturbances. The final diagnosis is derived from a holistic evaluation of these symptoms. However, this manual process is not only time-consuming and resource-intensive but also susceptible to the subjective judgment of individual clinicians, leading to potential inconsistencies.
To mitigate these limitations, automated depression diagnosis has emerged as a promising research avenue (Wei et al., 2022b; Yuan et al., 2024; Muzammel et al., 2021). Early approaches successfully analyzed dialogue semantics and extracted affective cues from speech and facial expressions, achieving high diagnostic efficiency and accuracy (Chen et al., 2024; Niu et al., 2020; Zhang et al., 2024; Xie and Peng, 2025). Yet, a critical flaw persists in these methods: their ”black-box” nature. While they can predict a depression score, they fail to provide a rationale for their predictions. This lack of explainability is a major barrier to clinical adoption, as practitioners are hesitant to trust a diagnostic tool without a clear, verifiable reasoning process.
The advent of Large Language Models (LLMs) (Yang et al., 2024; Touvron et al., 2023; Achiam et al., 2023; Sun et al., 2021; Zeng et al., 2022) offers a compelling solution to this explainability crisis. With their profound capabilities in understanding context (Zhao et al., 2023; Lei et al., 2023; Peng et al., 2023; Huang et al., 2024), performing logical reasoning (Hsieh et al., 2023; Wei et al., 2022a; Kojima et al., 2022), and generating coherent text (Raffel et al., 2020; Yang et al., 2024), LLMs can not only assess a patient’s depression level from interview transcripts but also articulate a detailed, evidence-based rationale for their conclusion. This ability to mirror the clinician’s process of ”scoring with explanation” promises to significantly enhance the trustworthiness and clinical utility of automated systems.
However, a purely text-based approach is fundamentally incomplete. Depression manifests not only in the content of a patient’s speech but also in their non-verbal behaviors. Cues such as vocal prosody (e.g., flat intonation) and facial expressions (e.g., lack of affect) are indispensable diagnostic indicators (Liu et al., 2023; De Melo et al., 2020) that provide a more objective and holistic view of a patient’s mental state. While state-of-the-art multimodal LLMs (MLLMs) like Gemini and Claude 3 can process raw video data (Team et al., 2023; Anthropic, 2024; Bai et al., 2025), they are impractical for real-world clinical applications. Due to stringent patient privacy regulations and ethical considerations, depression datasets almost never provide raw videos, but rather pre-extracted, anonymized features. This data modality gap renders existing end-to-end MLLMs unusable, presenting a significant technical challenge: how can we equip an LLMs with multimodal understanding capabilities when direct access to the original modalities is prohibited?
Furthermore, deploying frontier LLMs in clinical settings is often infeasible due to their massive computational footprint and operational costs. A practical solution demands the use of smaller, more efficient models. This introduces a second, equally critical challenge: how can a small-scale LLMs achieve the sophisticated reasoning and explanation-generation abilities typically associated with its large-scale counterparts, especially in a specialized domain like clinical psychology?
To address these intertwined challenges of modality alignment under privacy constraints and reasoning distillation for model practicality, we propose MLlm-DR, a novel Multimodal Large Language Model for explainable Depression Rationale generation. As illustrated in Figure 1, our framework introduces a solution that is both effective and practical. At its core, MLlm-DR is built upon a small-scale LLMs to ensure deployment feasibility. To imbue it with expert-level reasoning, we introduce a knowledge distillation strategy: we leverage a powerful, large-scale LLMs to generate high-quality diagnostic rationales from interview transcripts, creating a rich, explainable training corpus. Fine-tuning our small model on this distilled knowledge endows it with robust, clinically-aligned reasoning capabilities.
@@
To tackle the multimodal challenge, we design a novel and lightweight projection module, the Lightweight Query-based Transformer (LQ-former). Instead of processing raw data, LQ-former operates on pre-extracted audio and visual features, respecting data privacy. It employs a small set of learnable query vectors to dynamically probe the feature sequences and distill the most salient depression-related signals. These signals are then projected into a compact sequence of embeddings, effectively translating non-verbal cues into a representation the LLMs can process. We employ a principled, two-stage training strategy. First, we train only the LQ-former, teaching it to align multimodal features with the frozen LLM’s latent space. This decouples modality alignment from task-specific learning.
Second, with the trained LQ-former frozen, we fine-tune the LLMs using a synergistic multi-task objective. This objective jointly optimizes a cross-entropy loss for rationale generation and a Mean Squared Error (MSE) loss for the final score prediction. This dual-task optimization creates a powerful feedback loop: it compels the model to generate explanations that are not only textually coherent but also quantitatively consistent with the predicted depression score. This reciprocal constraint acts as a form of regularization, ensuring the generated rationales are more accurate and clinically grounded, thereby enhancing the model’s overall diagnostic fidelity.
Our contributions are summarized as follows:
  • We propose MLlm-DR, the first framework, to our knowledge, that enables a small-scale LLMs to perform explainable, multimodal depression diagnosis, addressing the key practical challenges of model efficiency and restricted data access.
  • We introduce a knowledge distillation strategy that effectively transfers the complex reasoning capabilities of a large-scale LLMs to a compact, efficient model, enabling it to generate high-quality, clinically-aligned diagnostic rationales.
  • We design a novel, lightweight projection module, LQ-former, that efficiently aligns pre-extracted audio-visual features with the LLM’s latent space, enabling effective multimodal understanding without requiring access to raw data.
  • Extensive experiments on the CMDC and E-DAIC-WOZ datasets demonstrate that our method achieves state-of-the-art performance, validating the effectiveness of our proposed framework in generating accurate and interpretable diagnostic results.
@@ 2.1. Multimodal Depression Recognition
Recently, multimodal fusion methods have made valuable progress in many depression recognition tasks. For instance, Wei et al. (Wei et al., 2022b) design independent attention fusion modules for each PHQ-8 sub-score to extract multimodal features relevant to specific sub-scores, generating individual sub-scores and ultimately aggregating them for overall depression evaluation. Yuan et al. (Yuan et al., 2024) introduce multi-order factor decomposition to extract features from single modalities and their cross-modal combinations, significantly enhancing the representational capacity of multimodal learning and improving model interpretability through a dynamic weighting mechanism. Jung et al. (Jung et al., 2024) explicitly model the hierarchical structure of interview questions (primary questions and follow-ups), simulating the diagnostic logic of clinicians while leveraging attention mechanisms to identify critical questions and features, enabling more precise depression detection. However, these methods rely on neural networks to directly predict depression scores without providing the corresponding rationale, which reveals a significant limitation.
@@ 2.2. LLMs-based Dialogue Understanding
Recently, large language models (LLMs) have made significant progress in dialogue understanding tasks. For instance, Li et al. (Li et al., 2024) demonstrate the potential of LLMs in Dialogue Relation Extraction (DRE) tasks, showing their superior ability to capture long-span and multi-turn dialogue information, outperforming traditional sequence-based and graph-based methods. Lei et al. (Lei et al., 2023) investigate the use of LLMs for Emotion Recognition in Conversations (ERC). They pretrain the model on a speaker identification task to capture emotional expression characteristics of different roles in dialogues and fine-tune it through multi-task learning by integrating ERC and emotion influence prediction tasks, enhancing performance. Additionally, Huang et al. (Huang et al., 2024) propose the Emotion-Cause Reasoning Chain (ECR-Chain) framework, leveraging LLMs to analyze statements in dialogues that trigger target emotions and thereby predict causal emotion entailment (CEE). This framework incorporates cognitive appraisal theory to deeply explore the process of emotion generation, thus providing strong interpretability for reasoning outcomes. These studies collectively demonstrate the tremendous potential of LLMs in advancing dialogue understanding tasks.
@@ 2.3. Cross-Modal Semantic Alignment
To enable LLMs to process and understand multimodal inputs, researchers have proposed various cross-modal semantic alignment methods(Alayrac et al., 2022; Wang et al., 2024), achieving significant progress in multiple cross-modal language generation tasks, such as visual question answering and image captioning. For example, Li et al. (Li et al., 2023) propose a semantic alignment module (Q-Former), which utilizes a set of learnable query vectors to capture key information from visual features. Q-Former is trained through a vision-language matching task to align visual and textual features, enabling LLMs to generate textual descriptions based on visual input. To reduce reliance on paired image-text data, Jian et al. (Jian et al., 2024) training their model on textual data to optimize prompts that guide the generation of language from visual inputs. Visual features are subsequently mapped to these prompts, achieving alignment between vision and language. To enhance the model’s understanding of fine-grained visual information, Lu et al. (Lu et al., 2023) capture multi-level image features through three submodules: image tagging, object detection, and semantic segmentation. This enables the LLMs to process more granular visual information. Additionally, approaches (Cheng et al., 2024) employ simple linear networks (e.g., MLP) to map non-textual features into the textual embedding space, facilitating modality alignment. These studies provide effective methods to extend text-only LLMs to process multimodal data.
@@ 3.1. Data Processing
To enhance the performance of small-scale LLMs in depression assessment tasks and enable them to generate more logically consistent evaluation rationales, we utilize large-scale LLMs (Achiam et al., 2023; Touvron et al., 2023) to generate evaluation rationales based on dialogue content and construct training dataset. Specifically, we provide a text instruction requesting the large-scale LLMs to produce an evaluation score within a range of 0-3 along with corresponding rationales based on the dialogue content. In this text instruction, we report the actual evaluation score of the patient, ensuring consistency between the rationales generated by the LLMs and the actual score, as shown in the Table 1. The training dataset is then used to fine-tune MLlm-DR, aiming to transfer the reasoning capabilities of the large-scale LLMs to a smaller model for domain-specific tasks. In addition, we use HuBERT (Hsu et al., 2021) to extract audio features from raw speech data. For visual features, we utilize the deep representations provided in the dataset, which are extracted using OpenFace 2.0 (Baltrusaitis et al., 2018) or ResNet 50 (He et al., 2016).
@@
User Given the participant’s self-rating score of {label}, please evaluate the participant’s performance in {aspect} based on the dialogue content. The output format is as follows: Evaluation Reason: A concise and logical description. Each output must strictly follow this format, avoiding omissions or confusion.
@@ 3.2. Lightweight Query Module
The LQ-former is designed to help LLMs process and understand multimodal information. It consists of a Transformer (Vaswani, 2017) Decoder and a Projection Module based on a fully connected network. The decoder takes a set of learnable query vectors $Q$ as input. In each decoder layer, the queries first interact with each other via self-attention, and then attend to the non-textual modality features (such as audio and visual embeddings) via cross-attention. In this way, the decoder encourages different queries to focus on different regions of the modality sequence and aggregate depression-related cues, producing fixed-length feature representations $H$. $H$ are then projected to the same dimensionality as the LLM’s textual embeddings through a Projection layer, producing the output $O$, as shown in the Equation  1. In our task, we utilize two LQ-former modules to process audio and visual features separately. The resulting features are concatenated with the tokenized textual instruction embeddings and used as inputs to the LLMs. To indicate the positions of the inserted features, we employ two special tokens, <AudioHere>and <VideoHere>, as markers.
@@
where $Q\in{\mathbb{R}^{n\times{D}}}$ is the query vector, while $K=V\in{\mathbb{R}^{m\times{D}}}$ represents the sequence of non-textual modality features, $D$ is the embedding dimension, $n$ denotes a fixed sequence length, and $m$ represents the length of the non-textual feature sequence.
3.3. Multi-Head Attention Aggregation Network
To process the regression prediction task, we design a Multi-Head Attention Aggregation Network (MAA). MAA takes the final hidden states $X\in\mathbb{R}^{L\times D}$ of the LLMs as input. It first splits $X$ into $h$ subspaces corresponding to $h$ attention heads, resulting in a tensor $x^{\prime}\in\mathbb{R}^{L\times h\times d}$, Each head then computes token-level attention weights $\alpha_{i}\in\mathbb{R}^{L}$ using a shared fully connected layer. These weights are used to aggregate the token representations into a sentence-level embedding. The embedding is passed through a fully connected layer to output the predicted scores $S_{\text{reg}}$, which are used to calculate the regression loss, as shown in the Equation  2.
$S_{\text{reg}}$ $=\text{FC}\left(\sum_{i=1}^{h}\left(\alpha_{i}\cdot x_{i}^{\prime}\right)\right),$
where $L$ is the sequence length, $h$ is the number of attention heads, and $d=D/h$ is the dimension of each subspace, $W_{i}\in\mathbb{R}^{d\times 1}$ is a trainable parameter for each head.
@@ 3.4. Multi-Task Joint Fine-Tuning
Existing LLMs primarily focus on text generation tasks, and their performance in score prediction is often limited when the provided prompt lacks sufficient example samples. To enhance the model’s performance in scoring prediction tasks, we introduce a joint optimization strategy combining causal language modeling cross-entropy loss ($\mathcal{L}_{\text{lm}}$) and mean squared error loss ($\mathcal{L}_{\text{mse}}$), as shown in the Equation  3.
(3) $\mathcal{L}_{\text{lm}}$ $=-\frac{1}{T}\sum_{t=1}^{T}\log p_{\theta}(x_{t}\mid x_{<t}),$
$\mathcal{L}_{\text{joint}}$ $=\mathcal{L}_{\text{lm}}+\mathcal{L}_{\text{mse}},$
where $\mathcal{L}_{\text{mse}}$ is the mean squared error between the predicted score and the ground-truth score; $\mathcal{L}_{\text{lm}}$ is the autoregressive language modeling loss, i.e., the token at position $t$ ($x_{t}$) is predicted only from the preceding tokens $x_{<t}$. $T$ is the sequence length, $p_{\theta}(x_{t}\mid x_{<t})$ is the model’s conditional probability of predicting $x_{t}$ given its context.